OpenStudy (anonymous):

what is the solution of the system? use substitution. 2x-y=-7, 4x-y=-4

OpenStudy (anonymous):

Rearrange the terms for either equation... Let's say the second equation. -y = -4x - 4 Substitute that for the first equation and solve for x... 2x - 4x - 4 = -7 -2x - 4 = -7 -2x = -3 x = 3/2 Finally... 2(3/2) - y = -7 3 - y = -7 -y = -10 y = 10 Hence, the solution is (3/2, 10) in terms of (x,y) I hope this helps!